# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@ 3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)
Table 13. Identified deployment challenge and solution for containerized applications in multi-cloud environment
@@
(Almeida et al., 2014) Performance Testing Deployment Standardized Benchmarking and Automation
(Verdugo et al., 2017) Deployment Validation Testing Process Management System
(Jha et al., 2019) Cost-effective Deployment Containerization and Microservice Architecture
(Pandey et al., 2022) Resource Management Deployment User Preference-Based Multi-Cloud Resource Brokering
(Bhamare et al., 2015) NFV Site Deployment Site Selection and VF Allocation Algorithms
(Serhiienko and Spillner, 2018) Cloud Management Platform Evaluation Standardized Output Formats and Evaluation Criteria
(Dimitrijevic et al., 2019) Standardized Requirement Deployment Application-Level Resource Managers for Multi-Cloud
(Baresi et al., 2016) Heterogeneous Node Deployment Technology-Independent Multi-Level Adaptation
(Casola et al., 2017) Deployment Plan Generation Cloud Services Selection and Deployment Planning
(Tricomi et al., 2017) TOSCA-based Deployment Brokered Multi-Cloud Deployment with TOSCA
(Sitaram et al., 2018) Standardization Deployment Proxy Cloud Virtualization for OpenStack
(Barletta et al., 2024) Criticality Constraints in Kubernetes Deployment Kubernetes for Critical Deployments
(da Silva et al., 2023) Automation Tools Deployment Issue Accurate Cost and Resource Tracking
(Pandi et al., 2023) Multi-cloud Deployment Challenge Jenkins for Multi-cloud Automation
(Sofia et al., 2024) Decentralized App Deployment Issue CODECO for Edge-Cloud Deployment
(Neeraj et al., 2023) SLA Monitoring Deployment Complexity Blockchain for SLA Data Integrity
(Petrovski and Gusev, 2024) Container and Serverless Architecture Deployment Issue AWS ECS and Lambda Deployment
(Solayman and Qasha, 2023) Heterogeneous Environment Deployment Challenge Edge-cloud Orchestration Management
(Sofia et al., 2023) Dynamic Adaptation Deployment Issue Context-aware Resource Allocation Strategies
(Santos et al., 2023) Regional Infrastructure Deployment Consistency Issue AWS Wavelength for Edge Cloud Integration
(Roy and Kalafatis, 2023) Robotic App Deployment Challenge Hybrid Edge-Cloud Architecture
(Alonso et al., 2023a) Multi-cloud Deployment Consistency Problem Multi-cloud Deployment Efficiency Models
(Harzenetter et al., 2023) Cross-Technology Deployment Integrity Issue Workflow Automation for Cross-Technology Deployment
(Huang and Wang, 2023) Industrial Robot Control System Deployment Problem Seamless Control between Cloud and Robots
(Nagarajan et al., 2023) Efficient Task Deployment across Clouds Hybrid Approach for Resource Allocation
@@
(Ferry et al., 2014) Vendor Lock-in Prevention Deployment Cloud Modelling Framework
(Gao et al., 2017) Security Deployment in Multi-container Environment Cross-Container Isolation
(Rios et al., 2017) Security Deployment Security Control Specification and Deployment Framework
(Alaluna et al., 2017) Virtual Network Security Deployment Network Virtualization Platform for Multi-Cloud
(Wahab et al., 2016) Malicious Service Management Deployment Collusion-Resilient Trust Aggregation Technique
(Barati et al., 2023) Secure Personal Data Deployment Challenge Docker, Kubernetes, Blockchain for GDPR Compliance
(Fugkeaw, 2023) IAM Deployment across Cloud Platforms Blockchain for Decentralized IAM Deployment
(Wang et al., 2023) Non-intrusive Malware Detection Deployment DockerWatch for Non-intrusive Container Integration
@@
(Sabbioni et al., 2020) Multi-cloud PaaS Deployment Lightweight Proxies for PaaS Integration
(Bou Ghantous and Gill, 2021) IoT Application Deployment DRA Framework and CI Broker for Multi-Cloud Deployment
(Ramesh et al., 2023) Data Transfer and Service Compatibility FaaS and Storage for Multi-Cloud Deployment
(Nguyen and Chien, 2023) Heterogeneous Cloud and Edge Integration RBAM for Cloud-Edge Integration
(Tiwari and Sharma, 2023) Private Cloud Integration Problem CLI Tool for Serverless Setup
@@
(Slominski et al., 2015) Legacy Code Migration Deployment Reusable Architectural Pattern with Docker and Cloudant
(Wang et al., 2020) Tenant Performance Deployment Multi-Tenant and Multi-Instance Hybrid Deployment
(Trihinas et al., 2015) Monitoring Deployment Platform-Independent Monitoring System
(Struhár et al., 2024) Resource Contention and Performance Problem Linux Patch for Dynamic CPU Adaptation
(Horchulhack et al., 2024) QoS Degradation Detection Problem Kubernetes with Performance Monitoring Tools
(Marques et al., 2024) Monitoring Integration in Hybrid Cloud Issue Custom Scaling with Kubernetes
(Kataru et al., 2023) AWS Deployment Performance Consistency AWS Model Comparison for Performance
(Benmerar et al., 2023) Cloud-Edge Deployment Performance Issue AI-driven Orchestration for Dynamic Resource Management
@@
(Barna et al., 2017) Autonomic System Deployment Self-Tuning Performance Model and Autonomic Management
(Fadda et al., 2019) Microservice Deployment Metrics and Requirements Framework for Microservices
(Pop et al., 2016) Multi-cloud Application Configuration Deployment Runtime Environment with Object Store and Artifact Repository
(Quint and Kratzke, 2018) Elastic Container Platform Dependency Deployment Separation of Elastic Platform and Cloud Application
(Hossain et al., 2023) Microservices Deployment in Edge Environments Docker, CI/CD, Service Mesh for Microservices
@@
(Carpio et al., 2020) Inter-edge Bandwidth Deployment Modular Edge Cloud Computing Architecture
(Shi et al., 2020) Network Latency Deployment VM Selection for Composite Applications
(Dreibholz et al., 2019) Latency Reduction Deployment Edge-Cellular Hybrid Infrastructure Provisioning
(Pellegrini et al., 2017) Vendor Lock-in Prevention Deployment Interoperability Layer Above Cloud Infrastructure
(Lee and Nam, 2023) Dynamic Network Policy Deployment Problem KUNERVA for Network Policy Automation
@@
(Pham et al., 2015) ACS Compatibility Deployment Scalable Multi-Cloud Deployment Framework
(Sheridan et al., 2017) Manual Intervention Risk Deployment Automated Security Measures at Deployment
(Wurster et al., 2018) Multi-provider Deployment TOSCA-Based Deployment Modeling Approach
(Challita et al., 2017) Cloud Provider Transition Deployment Semantic Interoperability in Multi-Clouds
(da Silva et al., 2023) Automation Tools Deployment Issue Accurate Cost and Resource Tracking
(Santos et al., 2023) Regional Infrastructure Deployment Consistency Issue AWS Wavelength for Edge Cloud Integration
@@
(Al-shammari and Alwan, 2018) Fog Computing Deployment Reorganized Container Images and Docker Deployment
(Sami and Mourad, 2020) Fog Device Deployment On-Demand Fog and Microservices Deployment
(Lorenzo et al., 2019) Fog Computing Container Deployment FogDocker for Container Deployment
(Altran et al., 2022) Distributed Compute Node Deployment Label-Based Scheduling Strategy
(Bélair et al., 2019) Kernel Security Deployment Enhanced Container Security
(Barletta et al., 2024) Criticality Constraints in Kubernetes Deployment Kubernetes for Critical Deployments
(Centofanti et al., 2024) Power Tool Compatibility Problem Docker, Kubernetes for Power Monitoring Integration
(Ramesh et al., 2023) Data Transfer and Service Compatibility FaaS and Storage for Multi-Cloud Deployment
@@
(Alonso et al., 2019) API Management Deployment DevOps Approach for Multi-Cloud Applications
(Gupta et al., 2021) Web Service Deployment Kubernetes-Based Containerized Deployment
(Buyya and Barreto, 2015) Complex Application Deployment Aneka Platform for Distributed Applications
@@
(Acquaviva et al., 2017) High-availability Deployment Middleware Support for High Availability in Multi-Cloud PaaS
(Kiss, 2018) Complex Workflow Deployment CloudSME Simulation Platform
@@
(Kritikos et al., 2019) Business Process Provisioning Deployment Multi-Cloud Service Orchestration Frameworks
(Roy and Kalafatis, 2023) Robotic App Deployment Challenge Hybrid Edge-Cloud Architecture
@@
(Sofia et al., 2023) Dynamic Adaptation Deployment Issue Context-aware Resource Allocation Strategies
Multi-cloud Deployment Coordination and Integration This category identifies and classifies the challenges and solutions related to coordination and integration during the multi-cloud deployment phase. Of the 121 reviewed studies, 5 addressed these challenges. One such challenge is Multi-cloud PaaS Deployment, which seeks to overcome high entry barriers for deploying a PaaS infrastructure across multiple clouds. The proposed solution is Lightweight Proxies for PaaS Integration, which allows for the seamless integration of different PaaS services through lightweight proxies. One other challenge includes IoT Application Deployment, which presents issues when deploying IoT applications on multi-cloud platforms. The solution for this challenge is the DRA Framework and CI Broker for Multi-Cloud Deployment, which ensure smooth deployment and resource coordination across multiple clouds. Additionally, Data Transfer and Service Compatibility issues are managed using FaaS and storage services, ensuring that data transfer and compatibility between different cloud providers are not problematic. RBAM supports heterogeneous cloud and edge integration, offering frictionless management and operational flexibility for cloud-edge configurations. Lastly, private cloud integration is facilitated by the CLI tools that enable easy integration and the establishment of serverless environments.
Monitoring and Scalability Challenges category identifies and classifies various issues and solutions related to monitoring, performance, and scalability in the context of multi-cloud container deployment. Of the 121 reviewed studies, 8 specifically addressed these issues. One challenge is Legacy Code Migration Deployment, which involves managing the migration of legacy code, tenant engine separation, and composition modeling during container deployment across multiple clouds. The proposed solution is a Reusable Architectural Pattern with Docker and Cloudant, which enables the reuse of an architectural pattern using Docker for containerization and Cloudant for the persistence layer. Another challenge is Tenant Performance Deployment, which focuses on managing tenant service performance and competition during multi-cloud containerized deployments. A proposed solution is the Multi-tenant and Multi-instance Hybrid Deployment scheme, based on container technology, which improves performance efficiency across tenants. A critical challenge is the monitoring of deployments, where existing tools have limitations in multi-cloud environment. The solution involves implementing JCatascopia, a platform-independent and interoperable monitoring system, which enhances the monitoring of multi-cloud setups and helps mitigate these challenges. Additional challenges include Resource Contention and QoS Degradation. The solutions include Dynamic CPU Adaptation through Linux patches, along with Kubernetes and performance monitoring tools. These measures ensure optimal resource usage and maintain performance consistency across cloud environments.
Network Connectivity and Hybrid Cloud Integration category classifies the challenges and solutions related to ensuring proper network connectivity in multi-cloud deployments and hybrid cloud integration. Out of the 121 studies reviewed, 5 specifically address these issues. One challenge, Inter-edge Bandwidth Deployment, deals with interconnecting distributed localities and managing bandwidth at the edge. The solution is a Modular Edge Cloud Computing Architecture, using containerization for better bandwidth management. A key challenge is Network Latency Deployment, which focuses on minimizing latency during multi-cloud deployments. The proposed solution is VM Selection for Composite Applications, optimizing virtual machine selection to reduce latency. Latency Reduction Deployment involves improving end-to-end performance through an Edge-Cellular Hybrid Infrastructure, combining edge and cellular networks. To address Vendor Lock-in, an Interoperability Layer Above Cloud Infrastructure enables smoother transitions between cloud providers. Finally, Dynamic Network Policy Deployment tackles the challenge of adapting network policies to container workloads, with the solution being Automation of Network Policies using tools like KUNERVA.
@@
[…]
3.8. Monitoring Challenges and Solution Framework (RQ4 and RQ5)
Table LABEL:tab:monitoring-solutions provides an extensive overview of the monitoring challenges and solution framework for containerized applications in multi-cloud environment. It encompasses diverse challenges related to monitoring containerized applications in the complex multi-cloud landscape. Each category in the figure targets specific monitoring aspects and offers solutions to address these challenges.

### v3

@@ 3.7 Deployment Challenges and Solution Framework (RQ4 and RQ5)
Multi-cloud Deployment Coordination and Integration: This category identifies and classifies the challenges and solutions related to coordination and integration during the multi-cloud deployment phase. Of the 121 reviewed studies, 5 addressed these challenges. One such challenge is Multi-cloud PaaS Deployment, which seeks to overcome high entry barriers for deploying a PaaS infrastructure across multiple clouds. The proposed solution is Lightweight Proxies for PaaS Integration, which allows for the seamless integration of different PaaS services through lightweight proxies. One other challenge includes IoT Application Deployment, which presents issues when deploying IoT applications on multi-cloud platforms. The solution for this challenge is the DRA Framework and CI Broker for Multi-Cloud Deployment, which ensure smooth deployment and resource coordination across multiple clouds. Additionally, Data Transfer and Service Compatibility issues are managed using FaaS and storage services, ensuring that data transfer and compatibility between different cloud providers are not problematic. RBAM supports heterogeneous cloud and edge integration, offering frictionless management and operational flexibility for cloud-edge configurations. Lastly, private cloud integration is facilitated by the CLI tools that enable easy integration and the establishment of serverless environments.
Monitoring and Scalability Challenges category identifies and classifies various issues and solutions related to monitoring, performance, and scalability in the context of multi-cloud container deployment. Of the 121 reviewed studies, 8 specifically addressed these issues. One challenge is Legacy Code Migration Deployment, which involves managing the migration of legacy code, tenant engine separation, and composition modeling during container deployment across multiple clouds. The proposed solution is a Reusable Architectural Pattern with Docker and Cloudant, which enables the reuse of an architectural pattern using Docker for containerization and Cloudant for the persistence layer. Another challenge is Tenant Performance Deployment, which focuses on managing tenant service performance and competition during multi-cloud containerized deployments.
A proposed solution is the Multi-tenant and Multi-instance Hybrid Deployment scheme, based on container technology, which improves performance efficiency across tenants. A critical challenge is the monitoring of deployments, where existing tools have limitations in multi-cloud environment. The solution involves implementing JCatascopia, a platform-independent and interoperable monitoring system, which enhances the monitoring of multi-cloud setups and helps mitigate these challenges. Additional challenges include Resource Contention and QoS Degradation. The solutions include Dynamic CPU Adaptation through Linux patches, along with Kubernetes and performance monitoring tools. These measures ensure optimal resource usage and maintain performance consistency across cloud environments.
Network Connectivity and Hybrid Cloud Integration category classifies the challenges and solutions related to ensuring proper network connectivity in multi-cloud deployments and hybrid cloud integration. Out of the 121 studies reviewed, 5 specifically address these issues. One challenge, Inter-edge Bandwidth Deployment, deals with interconnecting distributed localities and managing bandwidth at the edge. The solution is a Modular Edge Cloud Computing Architecture, using containerization for better bandwidth management. A key challenge is Network Latency Deployment, which focuses on minimizing latency during multi-cloud deployments. The proposed solution is VM Selection for Composite Applications, optimizing virtual machine selection to reduce latency. Latency Reduction Deployment involves improving end-to-end performance through an Edge-Cellular Hybrid Infrastructure, combining edge and cellular networks. To address Vendor Lock-in, an Interoperability Layer Above Cloud Infrastructure enables smoother transitions between cloud providers.
Table 13: Identified deployment challenges and solutions for containerized applications in multi-cloud environment
@@
      [95]        Performance Testing Deployment        Standardized Benchmarking and Automation
      [120]        Deployment Validation        Testing Process Management System
      [96]        Cost-effective Deployment        Containerization and Microservice Architecture
      [67]        Resource Management Deployment        User Preference-Based Multi-Cloud Resource Brokering
      [47]        NFV Site Deployment        Site Selection and VF Allocation Algorithms
      [45]        Cloud Management Platform Evaluation        Standardized Output Formats and Evaluation Criteria
      [61]        Standardized Requirement Deployment        Application-Level Resource Managers for Multi-Cloud
      [31]        Heterogeneous Node Deployment        Technology-Independent Multi-Level Adaptation
      [127]        Deployment Plan Generation        Cloud Services Selection and Deployment Planning
      [128]        TOSCA-based Deployment        Brokered Multi-Cloud Deployment with TOSCA
      [81]        Standardization Deployment        Proxy Cloud Virtualization for OpenStack
      [82]        Criticality Constraints in Kubernetes Deployment        Kubernetes for Critical Deployments
      [115]        Automation Tools Deployment Issue        Accurate Cost and Resource Tracking
      [117]        Multi-cloud Deployment Challenge        Jenkins for Multi-cloud Automation
      [138]        Decentralized App Deployment Issue        CODECO for Edge-Cloud Deployment
      [147]        SLA Monitoring Deployment Complexity        Blockchain for SLA Data Integrity
      [100]        Container and Serverless Architecture Deployment Issue        AWS ECS and Lambda Deployment
      [38]        Heterogeneous Environment Deployment Challenge        Edge-cloud Orchestration Management
      [118]        Dynamic Adaptation Deployment Issue        Context-aware Resource Allocation Strategies
      [125]        Regional Infrastructure Deployment Consistency Issue        AWS Wavelength for Edge Cloud Integration
      [46]        Robotic App Deployment Challenge        Hybrid Edge-Cloud Architecture
      [33]        Multi-cloud Deployment Consistency Problem        Multi-cloud Deployment Efficiency Models
      [141]        Cross-Technology Deployment Integrity Issue        Workflow Automation for Cross-Technology Deployment
      [49]        Industrial Robot Control System Deployment Problem        Seamless Control between Cloud and Robots
@@
      [89]        Vendor Lock-in Prevention Deployment        Cloud Modelling Framework
      [70]        Security Deployment in Multi-container Environment        Cross-Container Isolation
      [56]        Security Deployment        Security Control Specification and Deployment Framework
      [124]        Virtual Network Security Deployment        Network Virtualization Platform for Multi-Cloud
      [72]        Malicious Service Management Deployment        Collusion-Resilient Trust Aggregation Technique
      [135]        Secure Personal Data Deployment Challenge        Docker, Kubernetes, Blockchain for GDPR Compliance
      [144]        IAM Deployment across Cloud Platforms        Blockchain for Decentralized IAM Deployment
      [142]        Non-intrusive Malware Detection Deployment        DockerWatch for Non-intrusive Container Integration
@@
      [65]        Multi-cloud PaaS Deployment        Lightweight Proxies for PaaS Integration
      [90]        IoT Application Deployment        DRA Framework and CI Broker for Multi-Cloud Deployment
      [92]        Data Transfer and Service Compatibility        FaaS and Storage for Multi-Cloud Deployment
      [119]        Heterogeneous Cloud and Edge Integration        RBAM for Cloud-Edge Integration
      [131]        Private Cloud Integration Problem        CLI Tool for Serverless Setup
@@
      [44]        Legacy Code Migration Deployment        Reusable Architectural Pattern with Docker and Cloudant
      [66]        Tenant Performance Deployment        Multi-Tenant and Multi-Instance Hybrid Deployment
      [84]        Monitoring Deployment        Platform-Independent Monitoring System
      [137]        Resource Contention and Performance Problem        Linux Patch for Dynamic CPU Adaptation
      [87]        QoS Degradation Detection Problem        Kubernetes with Performance Monitoring Tools
      [88]        Monitoring Integration in Hybrid Cloud Issue        Custom Scaling with Kubernetes
      [134]        AWS Deployment Performance Consistency        AWS Model Comparison for Performance
      [104]        Cloud-Edge Deployment Performance Issue        AI-driven Orchestration for Dynamic Resource Management
@@
      [52]        Autonomic System Deployment        Self-Tuning Performance Model and Autonomic Management
      [85]        Microservice Deployment        Metrics and Requirements Framework for Microservices
      [32]        Multi-cloud Application Configuration Deployment        Runtime Environment with Object Store and Artifact Repository
      [50]        Elastic Container Platform Dependency Deployment        Separation of Elastic Platform and Cloud Application
      [37]        Microservices Deployment in Edge Environments        Docker, CI/CD, Service Mesh for Microservices
@@
      [30]        Inter-edge Bandwidth Deployment        Modular Edge Cloud Computing Architecture
      [97]        Network Latency Deployment        VM Selection for Composite Applications
      [122]        Latency Reduction Deployment        Edge-Cellular Hybrid Infrastructure Provisioning
      [73]        Vendor Lock-in Prevention Deployment        Interoperability Layer Above Cloud Infrastructure
      [145]        Dynamic Network Policy Deployment Problem        KUNERVA for Network Policy Automation
@@
      [86]        ACS Compatibility Deployment        Scalable Multi-Cloud Deployment Framework
      [55]        Manual Intervention Risk Deployment        Automated Security Measures at Deployment
      [140]        Multi-provider Deployment        TOSCA-Based Deployment Modeling Approach
      [94]        Cloud Provider Transition Deployment        Semantic Interoperability in Multi-Clouds
      [115]        Automation Tools Deployment Issue        Accurate Cost and Resource Tracking
      [125]        Regional Infrastructure Deployment Consistency Issue        AWS Wavelength for Edge Cloud Integration
@@
      [112]        Fog Computing Deployment        Reorganized Container Images and Docker Deployment
      [29]        Fog Device Deployment        On-Demand Fog and Microservices Deployment
      [113]        Fog Computing Container Deployment        FogDocker for Container Deployment
      [114]        Distributed Compute Node Deployment        Label-Based Scheduling Strategy
      [71]        Kernel Security Deployment        Enhanced Container Security
      [82]        Criticality Constraints in Kubernetes Deployment        Kubernetes for Critical Deployments
      [139]        Power Tool Compatibility Problem        Docker, Kubernetes for Power Monitoring Integration
      [92]        Data Transfer and Service Compatibility        FaaS and Storage for Multi-Cloud Deployment
@@
      [105]        API Management Deployment        DevOps Approach for Multi-Cloud Applications
      [103]        Web Service Deployment        Kubernetes-Based Containerized Deployment
      [98]        Complex Application Deployment        Aneka Platform for Distributed Applications
@@
      [146]        High-availability Deployment        Middleware Support for High Availability in Multi-Cloud PaaS
      [63]        Complex Workflow Deployment        CloudSME Simulation Platform
@@
      [76]        Business Process Provisioning Deployment        Multi-Cloud Service Orchestration Frameworks
      [46]        Robotic App Deployment Challenge        Hybrid Edge-Cloud Architecture
@@
      [118]        Dynamic Adaptation Deployment Issue        Context-aware Resource Allocation Strategies
Finally, Dynamic Network Policy Deployment tackles the challenge of adapting network policies to container workloads, with the solution being Automation of Network Policies using tools like KUNERVA.
@@
Takeaway 10: Reviewed studies show that deployment challenges in containerized multi-cloud environments emerge from the tight coupling of container orchestration with heterogeneous cloud infrastructures. Solutions such as TOSCA-based modeling, cross-container isolation, and context-aware adaptation specifically address complexities introduced when containers span multiple clouds, highlighting the unique integration, security, and coordination demands identified through this SMS.
3.8 Monitoring Challenges and Solution Framework (RQ4 and RQ5)
Table LABEL:tab:monitoring-solutions presents a categorized overview of monitoring challenges and corresponding solutions for containerized applications in multi-cloud environments. The framework captures a range of challenges related to performance variability, resource utilization, observability, and integration. Each of the six categories highlights a specific aspect of monitoring and summarizes relevant solutions proposed in the literature.